# Tailoring anisotropy in kirigami metamaterial skins with pop-up folding hinges

### v1

@@ Abstract
Kirigami metamaterial sheets and tubes, owing to their capacity to undergo large elastic deformations while developing three-dimensional surface textures, have enormous potential as skins for soft robots. Here, we propose to use kirigami skins with folding hinges in this same context. These recently-introduced kirigami feature counter-rotating panels connected by pop-up folding hinges. So far, researchers have only explored auxetic and highly-symmetric versions of such patterns. Yet, some of these attributes have to be relaxed in order to explore their full potential as robotic skins. Thus, we parameterize these patterns and relax symmetry constraints, with the goal of using this same platform to obtain a wide range of shape-morphing behaviors. We use a combination of: matrix analysis tools and analytical kinematic formulas to thoroughly explore the vast design space that ensues; experiments for validation purposes; and numerical simulations to explore the mechanics of selected planar and tubular patterns. We demonstrate that it is possible to tailor parameters to obtain skins that globally expand or contract due to axial elongation, and that present asymmetric pop-ups that can yield anisotropic friction — the most desired attribute for one-way locomotion of soft robots.
@@
Kirigami , Metamaterials , Shape morphing , Asymmetry , Soft robotics
@@ 1 Introduction
In recent decades, mechanical metamaterials have gained increasing attention due to their ability to manifest mechanical properties beyond conventional materials, such as negative Poisson’s ratio (auxeticity) [1, 2] and negative stiffness [3, 4]. These properties stem more from the geometry of the unit cells than from the material they are made of [5], and often emerge from the collective behavior of such unit cells. By altering the geometry of these building blocks, designers can completely alter the behavior of metamaterials. Of particular interest to us are metamaterials designed to undergo extreme shape changes and that can morph between distinct flat shapes [1, 6], from flat to 3D [7, 8] or between distinct 3D shapes [9, 10, 11, 12] when subjected to mechanical or non-mechanical stimuli. Such shape-morphing metamaterials have potential applications across scales, from nano- to micro-scale flexible electronic devices [13, 14], to tabletop-scale soft robots [15, 16, 17] and to large-scale morphing structures [18, 19, 20].
@@
Among shape-morphing metamaterials, particularly prominent are those that are based on origami and kirigami design principles [7, 21, 22, 23]. These involve thin sheets with folds, folds and cuts, or cuts alone. In origami and kirigami with folds and cuts, the presence of such discontinuities allow for the appearance of mechanism modes of deformation (referred to simply as mechanisms in the following) that manifest as large, low-energy relative rotations between panels [6, 24, 25]. In cut-only kirigami, the extreme morphing capacity can also stem from large relative rotations between panels [26, 27], or can be ascribed to localized [28, 29, 30, 31] or global [8, 32] buckling.
Due to their extreme morphing capacity, kirigami metamaterials have found applications in various areas of soft robotics, from grasping [33, 34] to actuation [35] and locomotion [36, 37, 38, 39]. Concentrating on crawling locomotion, kirigami sheets arranged in tubular geometries around a linear actuator have been used to create snake-inspired robots [36, 40]. This concept is shown in Fig. 1A, where we also illustrate some of the cut patterns that can be used in this context, inspired by Ref. [40].
@@
When the actuator pulls on the tubular skin, localized buckling mechanics cause the appearance of asymmetric pop-ups along the axis direction, i.e., the pop-ups are smooth if looked at from the robot tip and present cavities when looked at from the tail. This pop-up asymmetry can be leveraged to produce anisotropic friction — which is essential to control the direction of locomotion. Recent developments in this context have revolved around creating skins for reconfigurable steerability, whereby the same skin can provide anisotropic friction in different directions depending on how it is actuated [41]. Other works have explored the applicability of similar morphing principles to burrowing robots, designed to burrow underground following forms akin to those of worm-like organisms [42, 43].
@@
When designing kirigami skins for soft crawling robots, researchers have not been very concerned with the radial behavior of the skin, which typically contracts when expanded due to its positive Poisson’s ratio. Yet, for underground burrowing and other potential applications like pipe inspection, the radial behavior of the skin is very important, since the robot will have to exert pressure on the surrounding medium to propel itself [44, 45, 46, 47]. In turn, this would require skins with negative Poisson’s ratio, that radially expand when pulled axially. Yet, kirigami architectures that simultaneously present negative Poisson’s ratio and asymmetric pop-ups that can yield anisotropic friction, producing qualitative results as shown in Fig. 1B, are not common. One promising pattern in this sense was recently introduced by Tang et al. [37] and is shown in Fig. 1C. This is a variation of the rotating squares, planar kirigmi introduced by Grima and Evans [1], where the ideal point-like hinges are replaced with initially-flat folding hinges that pop up as the unit cell expands. This same geometry was reprised to study its applicability to elastic wave control [48], and its out-of-plane folding kinematics [49]. Yet, the highly symmetric versions of this pattern studied so far lack one of the fundamental attributes we are looking for: asymmetric pop-ups along the actuation direction.
@@
In this work, we start from the above-mentioned kirigami pattern with folding hinges and rationally explore its design space with the goal of identifying how to tune its parameters to tailor the effective Poisson’s ratio and produce asymmetric pop-ups. To do so, we introduce a generic parameterization of the unit cell and explore the relaxation of symmetry constraints within it. Our analysis involves: i) a multi-step study of the mechanisms of inextensional deformations in these patterns, done by using linear matrix structural analysis techniques [50, 26]; ii) a nonlinear kinematic analysis based on closed-form formulas of the unit cell’s deformation; iii) simple yet effective experiments on laser-cut unit cells to validate the kinematics; and iv) finite element (FE) simulations to probe the mechanics involved in the actuation of such patterns. While most of the study involves planar geometries, under the assumption that findings in this setting can be extrapolated to tubular patterns, we resort to simulations to model the behavior of a few tubular shells. Our study illustrates that it is indeed possible to find combinations of parameters that yield all the desired attributes. At the same time, our study provides comprehensive information on the morphing behaviors that can be achieved by kirigami with folding hinges, thus opening the way for others to use such patterns in soft robotics or other applications.
@@
This article is organized as follows. In Section 2, we introduce details on the unit cells and their parameters. In Section 3, we identify which combinations of parameters yield patterns with mechanisms, thus reducing the design space to a more manageable subset. In Section 4, we perform kinematic studies on the reduced design space and identify patterns with desired behaviors. In Section 5, we compare kinematic, numerical and experimental results for some representative patterns. Numerical results on tubular versions of some of our patterns are shown in Section 6. Concluding remarks are given in Section 7.
@@ 2 Unit cell geometry and design parameters
We begin by illustrating the kirigami unit with folding hinges by Tang et al. [37], and then introduce the necessary modifications to achieve the desired attributes of tailorable Poisson’s ratio and asymmetric pop-ups. Tang et al.’s unit cell is shown in Fig. 2A.
Fig. 2: Families of unit cell geometries and design parameters. (A) Unit cell geometry as introduced by Tang et al. [37], in which the only design parameters are representative lengths $L_{x}$ and $L_{y}$, fold length $\ell$ and fold angle $\lambda$; here the thicker lines represent cuts/slits. This geometry displays two mirror symmetry axes and, when actuated, evolves as shown in Fig. 1C. (B) Symmetric variation of the unit cell in (A), where we replace slits with diamond-shaped cuts with half-diagonals $w_{x}$ and $w_{y}$, and we allow for parameter variations across pop-ups ($\lambda_{i}$ and $\ell_{i}$, with $i=1,\ldots 4$, can be chosen independently), while maintaining mirror symmetries. (C) Asymmetric version of the unit cell in (B), where the pop-up labeling has changed as to potentially break mirror symmetries.
@@
It is composed of four octagonal panels (simply named panels in the following), each one inscribed within a square of size $L_{x}$ by $L_{y}$. It is assumed that the unit needs to be tiled periodically in space to create a larger pattern; thus, lattice vectors have magnitude of $2L_{x}$ and $2L_{y}$ along the horizontal and vertical directions, respectively. Each panel is connected to parts of its neighbors via folding hinges, here represented as lighter-gray, right triangles whose dimensions can all be obtained from their length $\ell$ and angle $\lambda$. Slits between panels, here shown as thicker lines, guarantee that the unit cell deforms according to the desired mechanism — labeled as global mechanism in the following, featuring in-plane counter-rotations of the panels and pop-up of the hinges, with all parts remaining rigid and deformations being concentrated at ideal fold lines as illustrated in Fig. 1C. This kirigami geometry possesses three distinct traits: i) it is symmetric about vertical and horizontal axes passing through the center of the cell; ii) it has negative Poisson’s ratio, just like the original rotating-squares geometry it takes inspiration from [1]; and iii) it features identical folding hinges that all pop up of the same amount.
@@
The first modification we make to this unit cell is to depart from line-like slits and replace them with diamond-shaped cuts. This echoes modifications made in the past on the rotating-squares pattern, that allowed to tune its effective Poisson’s ratio across a wide range of positive and negative values [8, 51]. In particular, as shown in Figs. 2B,C, each slit is either defined by its horizontal half-diagonal, $w_{x}$, or by its vertical one, $w_{y}$; its other dimension follows from its nearest neighboring cuts and from the folding hinges, as explained later. We here choose to define the central slit of the cell by $w_{y}$, its closest neighbors to the left, right, top and bottom by $w_{x}$, and the diagonally closest ones again by $w_{y}$ and so on.
@@
The second modification we apply is to relax the constraint that all folding hinges be identical. There are two ways we can do this. One way entails considering all hinges in the top-right quadrant of each unit cell different from each other, and to apply mirror symmetry about the Cartesian axes $x$ and $y$, to obtain a pattern as shown in Fig. 2B — labeled as symmetric in the following. Here and in the following, we color code hinges according to their type. We then define a hinge length parameter $\ell_{i}$ and an angle parameter $\lambda_{i}$ to characterize each hinge type, with $i=1,\ldots 4$. As a side note, hinge lengths $\ell_{i}$ are essential to define the undefined dimensions of the diamond-shaped cuts. For example, considering the central slit of the cell in Fig. 2B, we prescribe its vertical half-diagonal $w_{y}$; its horizontal half-diagonal, here identical for positive and negative $x$ due to symmetry, will be equal to $L_{x}-\ell_{4}$. It is useful to point out that patterns as in Fig. 2B are symmetric for all parameter choices.
@@
The third and final modification we introduce involves relaxing symmetry constraints about the $x$ and $y$ axes. We again consider the top-right quadrant of the cell and assign different parameters to all of its folding hinges. Then, we assign parameters to the hinges of the neighboring panels following these rules: i) each hinge needs to have the same properties in all its portions; ii) hinges that face each other across slits have to be of different types; iii) hinges have to be assigned so that periodicity is preserved while keeping the unit cell size to 2 $\times$ 2 panels. What results from this process are designs as the one in Fig. 2C — called asymmetric in the following — where any symmetry about the $x$ and $y$ axes is lost for most parameter choices, while their tileability in space to produce periodic patterns is preserved. These designs are the only ones that hold promise in the context of robotic skins with frictional anisotropy: if, for example, we consider the geometry in Fig. 2C and imagine using it for a robot that expands along the $y$ axis, then the fact that pop-ups of type 1 are larger than those of type 3 will likely cause the robot to move more easily along $+y$ than along $-y$, since larger pop-ups will generally yield higher friction. We spend the rest of the article to turn these qualitative statements into quantitative arguments, and to understand how to choose unit cell parameters to achieve a desired performance in terms of global and local morphing attributes.
@@
Prior to beginning our analysis, a few remarks on the design parameters are needed. First, throughout this work, we set the panel dimensions equal to each other: $L_{x}=L_{y}$. The characteristic dimensions of the cuts and the folding hinge lengths are normalized with respect to the panel dimension along the direction they are aligned with: $\bar{w}_{x}=w_{x}/L_{x}$, $\bar{w}_{y}=w_{y}/L_{y}$, ¯$\bar{\ell}_{1}=\ell_{1}/L_{y}$, ¯$\bar{\ell}_{2}=\ell_{2}/L_{x}$, ¯$\bar{\ell}_{3}=\ell_{3}/L_{y}$ and ¯$\bar{\ell}_{4}=\ell_{4}/L_{x}$.
3 Mechanism analysis and design space restriction
We now try to make some initial sense of the vast design space introduced in Section 2, with the goal of understanding which combinations of parameters allow us to preserve the global mechanism. To do so, we resort to linear matrix structural analysis arguments and, specifically, we adopt the framework of Pellegrino and Calladine [50], which allows to determine mechanisms and states of self stress of truss assemblies, and whether the mechanisms are infinitesimal or finite in nature. As we will see, the presence of the folding hinges leads to additional challenges.
The first step in adopting this formalism is to transform our panel-based unit cells into trusses. One such transformation between a symmetric geometry and its truss analog is shown in Figs. 3A,B.
To model panels as rigid, it is not enough to place bars along their outer boundaries, as this would yield highly floppy constructs. There are many ways to triangulate octagons and making them rigid; we choose to add a node to the center of each panel and connect all octagon nodes to the corresponding center node, as shown in Fig. 3B. Overall, the truss analog to our cell has $j=32$ joints and $b=64$ bars. In this truss structure, nodes can move out of plane; thus, the number of degrees of freedom per node is $n=3$. However, since we are interested in its planar morphing behavior (octagonal panels remain coplanar), we constrain all nodes to remain in the plane, apart from those at the tip of the pop-ups, shown in red in Fig. 3B. The planar boundary conditions we subject the unit to are shown in the same figure as pin and roller symbols. Importantly, these are chosen to be consistent with a mechanism motion that transforms an overall-rectangular cell to another rectangular shape — a constraint that needs to be satisfied by a unit within a periodic pattern.
The main steps of the Pellegrino-Calladine analysis are summarized in A. The 12 mechanisms of the truss in Fig. 3B are shown in Fig. 3C; by studying their interaction with the states of self stress, we determine that all these mechanisms are finite in nature. All of them are characterized by motions that are solely concentrated at the pop-ups, and there are as many mechanisms as the number of pop-ups. The fact that they are finite in nature indicates that these pop-ups are truly floppy and that, while they provide intriguing functionality to the metamaterial and affect its mechanics, they don’t affect the motion of the octagonal panels. Notably, the desired global mechanism is missing from this set. We have a conjecture for its absence: the desired mechanism can only manifest after the hinges pop-up; thus, there is a hierarchy of mechanisms in this system, that linear analysis tools cannot identify. We therefore perform a second analysis, in which we remove the hinges and construct the hinge-less planar truss in Fig. 3D, where panels are connected at single points. We subject this structure to the same boundary conditions and study its mechanisms. We find that it has a finite mechanism consistent with the expected deformation of the system.
Fig. 4: Classification of asymmetric geometries based on their mechanism response, and on whether they display finite, infinitesimal or no global mechanism. (A) Colormap obtained by setting slit openings $\bar{w}_{x}=\bar{w}_{y}=0$, hinge angles $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$, and by varying hinge lengths ¯$\bar{\ell}_{1}$, ¯$\bar{\ell}_{2}$, ¯$\bar{\ell}_{3}$ and ¯$\bar{\ell}_{4}$. Each box in the plot corresponds to a combination of ¯$\bar{\ell}_{1}$ and ¯$\bar{\ell}_{2}$; in turn, each square within each box corresponds to a combination of ¯$\bar{\ell}_{3}$ and ¯$\bar{\ell}_{4}$, as shown in detail (A-1), which is an enlarged view of a single box of (A). (B) Colormap similar to the one in (A), where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ vary and we fix $\bar{w}_{x}=\bar{w}_{y}=0$ and ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$. (B-1) enlarged view of a box of (B). (C-1)–(C-4) Some of the geometries in our design space. In all of them, the dashed line connects the rotation points, and is akin to a slit in classical rotating squares kirigami. (C-1) Design that has no mechanism, from (A), with ¯$\bar{\ell}_{1}=0.40$, ¯$\bar{\ell}_{2}=0.25$, ¯$\bar{\ell}_{3}=0.34$ and ¯$\bar{\ell}_{4}=0.10$. (C-2) Design that has an infinitesimal mechanism, from (A), with ¯$\bar{\ell}_{1}=0.18$, ¯$\bar{\ell}_{2}=0.28$, ¯$\bar{\ell}_{3}=0.22$ and ¯$\bar{\ell}_{4}=0.16$. (C-3) Design with finite mechanism, from (A), with ¯$\bar{\ell}_{1}=\bar{\ell}_{3}=0.40$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}=0.25$. (C-4) Design with finite mechanism, from (B), with $\lambda_{1}=\pi/4$, $\lambda_{2}=\pi/8$, $\lambda_{3}=\pi/16$, $\lambda_{4}=\pi/32$.
The results of this parametric analysis are shown in Fig. 4. In the first study, we set $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\pi/4$ and investigate the effects of the individual hinge lengths. To do so, we use a nested colormap plot where, for each combination of ¯$\bar{\ell}_{1}$ and ¯$\bar{\ell}_{2}$ in a desired range, we plot a colormap as a function of ¯$\bar{\ell}_{3}$ and ¯$\bar{\ell}_{4}$, as shown in Fig. 4A. A detail of a colormap for specific values of ¯$\bar{\ell}_{1}$ and ¯$\bar{\ell}_{2}$ is shown in inset A-1. Points of the colormap are marked dark red if that combination of parameters yields a global mechanism of finite nature, pink if it yields an infinitesimal mechanism and white if it does not feature a mechanism. The range of values chosen for the normalized hinge lengths (0.1–0.4) avoids panel overlaps in the pattern. The result of this analysis is striking: to have a finite mechanism, we need to always have that ¯$\bar{\ell}_{1}=\bar{\ell}_{3}$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}$. In other words, hinges across slits must have the same length. In the second study, we instead fix ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$ and we let the folding angles $\lambda_{i}$ vary. Again, we use a similar nested colormap to compactly present four-dimensional data; the results are shown in Fig. 4B and an inset of one of the individual colormaps is shown in Fig. 4B-1. The result is again striking: if all hinge lengths are identical, hinge angles can all be chosen independently of one another, as this does not affect the global mechanism. The implications of this analysis are very important for the rest of the article, and are summarized below:
  • 1.
    the fact that we cannot choose hinge lengths independently leads to a significant restriction of the design space;
  • 2.
    the fact that we can instead choose hinge angles independently opens the door towards designs with asymmetric pop-ups.
Further restrictions of the design space come from the fact that not all parameters can be chosen independently, as this could yield overlaps between panels.
To understand why hinge lengths cannot be chosen independently, we consider some of the geometries sampled form our design space, shown in Figs. 4C-1–C-4, and add a dash line connecting the points about which the panels rotate. We can see that in C-1 and C-2 — designs that have all different values of hinge length and that yield infinitesimal or no mechanisms — the line connecting the rotation points is not a parallelogram. On the other hand, in the C-3 and C-4 designs — that feature finite mechanisms and where hinge lengths are either identical or such that opposite hinges have equal length, and where the angles are either identical or all different — the dashed line forms a parallelogram; thus, this seems to be the ingredient needed to have a mechanism. This is not entirely surprising: the dashed line here is akin to a slit in a rotating squares metamaterial, and other studies have pointed out the necessity of having parallelogram slits to produce mechanisms in rotating squares and their variations [52, 53, 54, 51].

### v2

# Tailoring asymmetry for anisotropic friction in kirigami metamaterial skins with pop-up folding hinges

@@ Abstract
Kirigami metamaterial sheets and tubes, owing to their capacity to undergo large elastic deformations while developing three-dimensional surface textures, have enormous potential as skins for soft robots. Here, we propose to use kirigami skins with folding hinges in this same context. These recently-introduced kirigami feature counter-rotating panels connected by pop-up folding hinges. So far, researchers have only explored auxetic and highly-symmetric versions of such patterns. Yet, some of these attributes have to be relaxed in order to explore their full potential as robotic skins. Thus, we parameterize these patterns and relax symmetry constraints, with the goal of using this same platform to obtain a wide range of shape-morphing behaviors. We derive kinematic formulas to explore the vast symmetry-enabled design space. We then use numerical simulations and experiments to validate the kinematic predictions and to explore the morphing mechanics of tubular skins. Finally, via experiments, we provide preliminary evidence of the anisotropic friction enabled by patterns with asymmetric pop-ups. We therefore demonstrate that it is possible to tailor parameters in kirigami with folding hinges to obtain skins that globally expand or contract due to axial elongation, and that present asymmetric pop-ups that yield anisotropic friction — the most desired attribute for one-way locomotion of soft robots.
This article may be downloaded for personal use only. Any other use requires prior permission of the authors and Elsevier. This article appeared in: International Journal of Mechanical Sciences 296, 110258 (2025) and may be found at: https://doi.org/10.1016/j.ijmecsci.2025.110258
@@
Kirigami , Metamaterials , Soft Robotics , Asymmetry , Auxeticity , Anisotropic Friction
@@ 1 Introduction
In recent decades, mechanical metamaterials have gained increasing attention due to their ability to manifest mechanical properties beyond conventional materials, such as negative Poisson’s ratio (auxeticity) [1, 2], negative stiffness [3, 4], or quasi-zero-stiffness [6, 5]. These properties stem more from the geometry of the unit cells than from the material they are made of [7], and often emerge from the collective behavior of such unit cells. By altering the geometry of these building blocks, designers can completely alter the behavior of metamaterials. Of particular interest to us are metamaterials designed to undergo extreme shape changes and that can morph between distinct flat shapes [1, 8], from flat to three dimensional (3D) [9, 10] or between distinct 3D shapes [11, 12, 13, 14] when subjected to mechanical stimuli. Such shape-morphing metamaterials have potential applications across scales, from nano- to micro-scale flexible electronic devices [15, 16], to tabletop-scale soft robots [17, 18, 19] and to large-scale morphing structures [20, 21, 22]. Metamaterials can also be designed to exhibit multiple stable configurations with distinct shapes [23, 24, 25, 26, 27] — an attribute that can be leveraged for energy harvesting, sensing, and robotics applications. Beyond purely mechanical stimuli, researchers are also investigating the use of other triggers to induce shape changes in metamaterials, such as light, heat, and magnetic fields  [28, 29, 30, 31, 32, 33].
@@
Among shape-morphing metamaterials, particularly prominent are those that are based on origami and kirigami design principles [9, 34, 35, 36, 48]. These involve thin sheets with folds, folds and cuts, or cuts alone. In origami and kirigami with folds and cuts, the presence of such discontinuities allow for the appearance of mechanism modes of deformation (referred to simply as mechanisms in the following) that manifest as large, low-energy relative rotations between panels [8, 37, 38]. In cut-only kirigami, the extreme morphing capacity can also stem from large relative rotations between panels [39, 40], or can be ascribed to localized [41, 42, 43, 44] or global [10, 45] buckling. Such complex morphing attributes of origami and kirigami metamaterials have made them particularly appealing in bio-inspired engineering. On one hand, such metamaterials can be leveraged to replicate the morphing behavior of natural systems [47, 46]; on the other end, some natural systems owe their morphing behavior to origami and kirigami-like designs [49, 48].
Owing to their attributes and potential in bio-inspiration, origami and kirigami metamaterials are becoming prevalent in soft robotics. For example, robots inspired by origami principles can fold into compact forms for navigation through constrained environments and subsequently unfold to perform more complex tasks [48, 50, 51]. Origami crawlers and grippers are also extremely popular [52, 53, 18]. Due to their extreme morphing capacity, kirigami metamaterials have also found applications in various areas of soft robotics, from grasping [54, 55] to actuation [56] and locomotion [57, 58, 59, 60]. Locally-buckling kirigami sheets, in particular, are a great platform to achieve directional friction, which is essential for efficient crawling locomotion [57, 63, 61]. Specifically, when wrapping such sheets in tubular shapes and when pulling on them axially via a linear actuator, the tailored cuts of the kirigami can produce pop-ups that remain flush in one direction, enabling smooth low-friction movement, while engaging with the surface When displaced in the opposite direction, increasing friction and providing enhanced grip. The resulting directional friction facilitates directed locomotion, promoting forward movement while mitigating backward slippage, thus effectively mimicking the mechanics of crawling organisms such as snakes [57, 62]. This concept is shown in Fig. 1(a), where we also illustrate some of the cut patterns that can be used in this context, inspired by Ref. [62]. Recent developments in this context have revolved around creating skins for reconfigurable steerability, whereby the same skin can provide anisotropic friction in different directions depending on how it is actuated [63]. Other works have explored the applicability of similar morphing principles to burrowing robots, designed to burrow underground following forms akin to those of worm-like organisms [64, 65].
@@
When designing kirigami skins for soft crawling robots, researchers have not been very concerned with the radial behavior of the skin, which typically contracts when expanded due to its positive Poisson’s ratio. Yet, for underground burrowing and other potential applications like pipe inspection, the radial behavior of the skin is very important, since the robot will have to exert pressure on the surrounding medium to propel itself [66, 67, 68, 69]. In turn, this would require skins with negative Poisson’s ratio, that radially expand when pulled axially. Yet, kirigami architectures that simultaneously present negative Poisson’s ratio and asymmetric pop-ups that can yield anisotropic friction, producing qualitative results as shown in Fig. 1(b), are not common. One promising pattern in this sense was recently introduced by Tang et al. [58] and is shown in Fig. 1(c). This is a variation of the rotating squares, planar kirigmi introduced by Grima and Evans [1], where the ideal point-like hinges are replaced with initially-flat folding hinges that pop up as the unit cell expands. This same geometry was reprised to study its applicability to elastic wave control [70], and its out-of-plane folding kinematics [71]. Yet, the highly symmetric versions of this pattern studied so far lack one of the fundamental attributes we are looking for: asymmetric pop-ups along the actuation direction, as shown in Fig. 1(d).
@@
In this work, we start from the above-mentioned kirigami pattern with folding hinges and rationally explore its design space with the goal of identifying how to tune its parameters to tailor the effective Poisson’s ratio and produce asymmetric pop-ups. To do so, we introduce a generic parameterization of the unit cell and explore the relaxation of symmetry constraints within it. Our study begins with a kinematic analysis of the unit cell’s deformation, to probe the vast design space and understand how folding hinge parameters relate to in-plane kinematics; this analysis yields a few representative symmetric and asymmetric patterns which we subject to further analysis. We perform finite element (FE) simulations of unit cells and experiments aimed at validating the kinematic predictions. We then fabricate and test tubular skins and to understand their mechanics and the influence of the sheet’s thickness, comparing these results to numerical simulations. Finally, we perform experiments aimed at validating our claim on the anisotropic friction of patterns with asymmetric pop-ups. Our results illustrate that it is indeed possible to find combinations of parameters that yield all the desired attributes: tunable Poisson’s ratio, asymmetric pop-ups and, consequently, anisotropic friction. Additionally, they highlight the importance of choosing the correct parameters for the tubular skins behave as their planar counterparts. In summary, our study provides comprehensive information on the morphing behaviors that can be achieved by kirigami with folding hinges, thus opening the way for others to use such patterns in soft robotics or other applications.
@@
This article is organized as follows. In Section 2, we introduce details on the unit cells and their parameters. In Section 3, we perform kinematic studies to explore the vast parameter space of kirigami with folding hinges and identify patterns with desired behaviors. In Section 4, we compare kinematic, numerical and experimental results for some representative patterns. Numerical and experimental results on the morphing mechanics of tubular skins, and preliminary results on the anisotropic friction stemming from asymmetric patterns, are reported in Section 5. Concluding remarks are given in Section 6.
Fig. 2: Families of unit cell geometries and design parameters. (a) Unit cell geometry as introduced by Tang et al. [58], in which the only design parameters are representative lengths $L_{x}$ and $L_{y}$, fold length $\ell$ and fold angle $\lambda$; here the thicker lines represent cuts/slits. This geometry displays two mirror symmetry axes and, when actuated, evolves as shown in Fig. 1(c). (b) Symmetric variation of the unit cell in (a), where we replace slits with diamond-shaped cuts with half-diagonals $w_{x}$ and $w_{y}$, and we allow for parameter variations across pop-ups ($\lambda_{i}$ and $\ell_{i}$, with $i=1,\ldots 4$, can be chosen independently), while maintaining mirror symmetries. (c) Asymmetric version of the unit cell in (b), where the pop-up labeling has changed as to potentially break mirror symmetries.
@@ 2 Unit cell geometry and design parameters
The kirigami unit with folding hinges by Tang et al. [58] is shown in Fig. 2(a). It is composed of four octagonal panels, each one inscribed within a square of size $L_{x}$ by $L_{y}$. It is assumed that the unit needs to be tiled periodically in space to create a larger pattern; thus, lattice vectors have magnitude of $2L_{x}$ and $2L_{y}$ along the horizontal and vertical directions, respectively. Each panel is connected to parts of its neighbors via folding hinges, here represented as lighter-gray triangles whose dimensions can all be obtained from their length $\ell$ and angle $\lambda$. Slits between panels, here shown as thicker lines, guarantee that the unit cell deforms according to the desired mechanism, featuring in-plane counter-rotations of the panels and pop-up of the hinges, with all parts remaining rigid and deformations being concentrated at ideal fold lines as illustrated in Fig. 1(c). This kirigami geometry possesses three distinct traits: it is symmetric about vertical and horizontal axes passing through the center of the cell; it has negative Poisson’s ratio; and it features identical folding hinges that all pop up of the same amount.
@@
The first modification we introduce is to depart from line-like slits and replace them with diamond-shaped cuts. This echoes modifications made in the past to the rotating-squares pattern, that allowed to tune its effective Poisson’s ratio across a wide range of positive and negative values [10, 73]. As shown in Figs. 2(b), (c), each slit is either defined by its horizontal half-diagonal, $w_{x}$, or by its vertical one, $w_{y}$; its other dimension follows from its nearest neighboring cuts and from the folding hinges. We here choose to define the central slit of the cell by $w_{y}$, its closest neighbors to the left, right, top and bottom by $w_{x}$, and the diagonally closest ones again by $w_{y}$ and so on.
@@
The second modification we apply is to relax the constraint that all folding hinges are identical. There are two ways we can do this. One way entails considering all hinges in the top-right quadrant of each unit cell different from each other, and to apply mirror symmetry about the Cartesian axes $x$ and $y$, to obtain a pattern as shown in Fig. 2(b) — labeled as symmetric in the following. Here and in the following, we color code hinges according to their type. We then define a hinge length parameter $\ell_{i}$ and an angle parameter $\lambda_{i}$ to characterize each hinge type, with $i=1,\ldots 4$. It is useful to point out that patterns as in Fig. 2(b) are symmetric for all parameter choices.
@@
The third and final modification we introduce involves relaxing symmetry constraints about the $x$ and $y$ axes. We again consider the top-right quadrant of the cell and assign different parameters to all of its folding hinges. Then, we assign parameters to the hinges of the neighboring panels following these rules: each hinge needs to have the same properties in all its portions; hinges that face each other across slits have to be of different types; and hinges have to be assigned so that periodicity is preserved while keeping the unit cell size to 2 $\times$ 2 panels. What results from this process are designs as the one in Fig. 2(c) — called asymmetric in the following — where any symmetry about the $x$ and $y$ axes is lost for most parameter choices, while their “tileability” in space to produce periodic patterns is preserved. These designs are the only ones that hold promise in the context of robotic skins with frictional anisotropy: if, for example, we consider the geometry in Fig. 2(c) and imagine using it for a robot that expands along the $y$ axis, then the fact that pop-ups of type 1 are larger than those of type 3 will likely cause the robot to move differently along $+y$ and $-y$ directions.
@@
Throughout this work, we set the panel dimensions equal to each other: $L_{x}=L_{y}$. Additionally, the characteristic dimensions of the cuts and the folding hinge lengths are normalized with respect to the panel dimension along the direction they are aligned with: $\bar{w}_{x}=w_{x}/L_{x}$, $\bar{w}_{y}=w_{y}/L_{y}$, ¯$\bar{\ell}_{1}=\ell_{1}/L_{y}$, ¯$\bar{\ell}_{2}=\ell_{2}/L_{x}$, ¯$\bar{\ell}_{3}=\ell_{3}/L_{y}$ and ¯$\bar{\ell}_{4}=\ell_{4}/L_{x}$.